OpenStudy (shelby1290):

State the missing element of each ordered pair so that the ordered pair belongs to the solution set of the equation. a) x+y=9; -> (4,__), (__,1), (-3,__), (__,-5), (12,__) b) x-y=6; -> (9,__), (__,2), (5,__), (__,-3), (-8,__) c) 2x+y=15; -> (1,__), (__,3), (5,__), (__,-5), (7,__)

OpenStudy (jadeishere):

well, it's kind of just subtraction, for example x+y = 9 so if x = 4 then y =5

OpenStudy (shelby1290):

ok then how about for the second one (___,1) @Jadeishere

OpenStudy (jadeishere):

Well, try it

OpenStudy (shelby1290):

I'm not so sure of how I'm supposed to subtract it

OpenStudy (shelby1290):

Is there some sort of rule to it?

OpenStudy (shelby1290):

@Jadeishere ^^^

OpenStudy (jadeishere):

Sorry, anywho, no. It's just 9- the number given and you'll get your answer

OpenStudy (shelby1290):

kk so 9-1=8 soo...(8,1) alright what about with x-y=6

OpenStudy (jadeishere):

the same concept :)

OpenStudy (jadeishere):

Kind of! Since it is x - y = 6, it's going to be (for example) 9 - y = 6

OpenStudy (shelby1290):

Okay, I believe i understand that... and what about 2x+y=15

OpenStudy (jadeishere):

Again, i'll be using the first pair if 2x + y = 15 then substitute the x value into the equation 2(1) + y = 15 then 3 + y = 15, subtract 3 from 15, and y = 12 :) Think you can do the rest?

OpenStudy (shelby1290):

How'd you get 3 isnt y 13

OpenStudy (shelby1290):

2x1+13=15

OpenStudy (jadeishere):

Oh oops!! Good thing you caught me /).<

OpenStudy (shelby1290):

lol

OpenStudy (shelby1290):

wait, was i right for that one? y=13

OpenStudy (shelby1290):

im so confused

OpenStudy (jadeishere):

Yeah, you're right

OpenStudy (shelby1290):

what about (___,3)

OpenStudy (jadeishere):

just plug in the value 2x + (3) = 15 2x = 12 x = 6

OpenStudy (shelby1290):

okay thank you so much

OpenStudy (jadeishere):

Of course :)